A physics lecture room has a volume of 248 m3 .

Part A

For a pressure of 1.00 atm and a temperature of 26.0 ∘C, use the ideal-gas law to estimate the number of air molecules in the room. Assume all the air is N2.

Part B

Calculate the particle density-that is, the number of N2 molecules per cubic centimeter.

Express your answer to three significant figures and include the appropriate units.

Part C

Calculate the mass of the air in the room.

Express your answer to three significant figures and include the appropriate units.
Answer #1

A)

Moles, n = PV/RT

n = (1 x 248000)/(0.0821 x 299) = 10102.70 mol

Molecules = 10102.70 x 6.02 x 10^23 = 6.08 x 10^27

B)

Density = (6.08 x 10^27)/(248000000) = 2.45 x 10^19 molecules/cm^3

C)

Mass, m = 10102.70 x 28 = 2.83 x 10^5 kg

Comment in case any doubt please rate my answer ...